can u plz solve these questions it is of wavy curve method


1/(x-2)(3-x) <0


-2/(x+3)(4-x) is greater than or equal to 0.

Grade:11

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:Hello student, please find answer to your question
\frac{1}{(x-2)(3-x)} < 0
\frac{1}{(x-2)(x-3)} > 0
By wavy curve method, we have
x < 2, x > 3

\frac{-2}{(x+3)(4-x)} \geq 0
\frac{2}{(x+3)(x-4)} \geq 0
Using wavy curve method
x<-3, x > 4